OpenStudy (anonymous):

Calc AB Question I have an answer, I just need to know if my answer is correct. The function v(t) = t^2 - 5t is the velocity in m/sec of a particle moving along the x-axis, where t is measured in seconds. Write an expression for the displacement and the total distance traveled by the particle on the interval 0

OpenStudy (anonymous):

u just need to integrate this eq with limits 0 and 10

OpenStudy (anonymous):

My answer is \[s(t) = \int\limits_{0}^{10}(t^2-5t)dt\]

OpenStudy (anonymous):

All it says to do is write the expression, it doesn't say to solve.

OpenStudy (anonymous):

s(t) is displacement right?

ganeshie8 (ganeshie8):

yes it gives u s(10) - s(1)

ganeshie8 (ganeshie8):

so clearly its a displacement

ganeshie8 (ganeshie8):

*s(10) - s(0)

OpenStudy (anonymous):

Lol ok I'm just making sure.

ganeshie8 (ganeshie8):

for distance u need to take abs value of velocity function

OpenStudy (anonymous):

Yes I know. @sourwing taught me how to do distance.

ganeshie8 (ganeshie8):

distance : \(\large s(t) = -\int\limits_{0}^{5}(t^2-5t)dt + \int\limits_{5}^{10}(t^2-5t)dt \)

ganeshie8 (ganeshie8):

okay good :)

OpenStudy (anonymous):

s(t) isn't displacement, it's position. displacement = ∫ v(t) dt , from a to b

OpenStudy (anonymous):

I thought x(t) was position

ganeshie8 (ganeshie8):

distance : \(\large s(t) = \int |(t^2-5t)| dt \) displacement : \(\large s(t) = \int (t^2-5t)dt \)

ganeshie8 (ganeshie8):

for dispacement u can simply take take : s(10) - s(0)

OpenStudy (anonymous):

those x and s are just letters they pick to represent the position. Certainly you could have easily chosen p(t) to represent the position The reason I said s(t) isn't displacement is because it is position. if s(5) = 5, this only means the position is 5 unit away from 0. Displacement is change in position, so you will need two t value

ganeshie8 (ganeshie8):

^^

OpenStudy (anonymous):

Ohhhhhhhhh lol (it clicked)

OpenStudy (anonymous):

that's why displacement = ∫ v(t) dt, from a to b if you just do ∫v(t) dt without the limit of integration, then it is a position function

OpenStudy (anonymous):

Oh true

OpenStudy (anonymous):

Well thanks again @sourwing !! That's the last question for today haha